To estimate the company's WACC, Marshall Inc. recently hired you as a consultant. You have obtain...

  1. To estimate the company's WACC, Marshall Inc. recently hired you as a consultant. You have obtained the following information.

(1) The firm's existing noncallable bonds which mature in 40 years, have an 5.00% annual coupon, a par value of $1,000, and a market price of $950. You have done some research and estimate the cost of issuing additional debt would cost you similarly to the existing bonds.

(2) The company's current tax rate is 40%, but the tax rate is estimated to go up to 35% very soon.

(3) The projected future risk-free rate is 2.50%. The market return is predicted to be 7.50%. The stock's historical beta is 1.52, as some uncertainty resolved, it’s expected to decrease to 1.20.

(4) The target capital structure consists of 25% debt and the balance is common equity. While based on the book value, debt accounts for 10% and equity accounts for 90%.

The firm uses CAPM to estimate the cost of common stock, and it does not expect to issue any new shares.

What is its WACC given all above information?
Answer #1

Working 1: Calculation of expected cost of common stock using CAPM :

Cost of stock = Rf + Beta (Rm - Rf)

Here,

Rf (Risk free return) = 2.50% or 0.025

Rm (Market rate) = 7.50% or 0.075

Beta (Expected) = 1.20

Now, put the values into the formula,

Cost of stock = 0.025 + 1.20 * (0.075 - 0.025)

Cost of stock = 0.025 + 0.06

Cost of stock = 0.085 or 8.50%

Working 2 : Calculation of Cost of debt :

Here cost of debt is equals to YTM (yield to maturity) of bond.

YTM of bond= (Copoun + ((P - M) /n)) / ((P + M)/2)

P (Par value) = $1000

M (Market price) = $950

n (years) = 40 years

Copoun = Par value * rate = $1000 * 5% = $50

Now, put the values into formula,

YTM = ($50 + (($1000 - $950)/40)) / (($1000 + $950) / 2)

YTM = ($50 + $1.25) / $975

YTM = 0.0526 or 5.26%

So, cost of debt = 5.26% (Cost of debt = YTM)

Working 3 : Calculation of WACC (using target capital structure)

Cost of equity = 8.50% or 0.085

Cost of debt = 5.26% or 0.0526

Weight of debt = 25% or 0.25

Weight of equity = 75% or 0.75 (100 - 25 ie debt)

Tax rate (Estimated) = 35% or 0.35

Now,

WACC = (Weight of debt * Cost of debt * (1 - tax rate)) + (Weight of equity * Cost of equity)

WACC = (0.25 * 0.0526 *(1 - 0.35)) + (0.75 * 0.085)

WACC = (0.25 * 0.0342) + 0.0638

WACC = 0.0086 + 0.0638

WACC (as per target capital structure) = 0.0724 or 7.24%

Working 4 : WACC calculation as per book value:

Cost of equity = 8.50 % or 0.085

Cost of debt = 5.26% or 0.0526

Weight of debt = 10% or 0.10

Weight of equity = 90% or 0.90

Tax rate = 35% or 0.35

Now,

WACC = (Weight of debt * Cost of debt * (1 - tax rate)) + (Weight of equity * Cost of equity)

WACC = (0.10 * 0.0526 *(1 - 0.35)) + (0.90 * 0.085)

WACC = (0.10 * 0.0342) + 0.0765

WACC = 0.0034 + 0.0765

WACC(as per book value) = 0.0799 or 7.99%